An IEEE 802.3 MAC Frame has 8 octet addresses. A message may be unicast, multicast, or broadcast. If the LSB of the first octet of the destination address is 0, the address is unicast; otherwise it is multicast. A broadcast message has all 1s in its address. If destination addresses of three messages are 4B:6A:03:49:5C:F0:44:A6:05:E5:32:0B , and FF:FF:FF:FF:FF:FF, state their types.
Answer #1

Given destination address are represented in hexadecimal form. Each letter in hexadecimal form is converted into binary using 4 bit form.

Message1 : 4B:6A:03:49:5C:F0

The LSB of the first octet of the destination is F0 = 11110000 ( not equal to zero)

According to given rule the LSB of the first octet of the destination address is not equal to zero then it is multicast

Message2 : FF:FF:FF:FF:FF:FF

The LSB of the first octet of the destination is FF = 11111111 ( all bits are equal to 1's)

According to given rule the LSB of the first octet of the destination address have all 1's then it is broadcast

Message2 : 44:A6:05:E5:32:0B

The LSB of the first octet of the destination is 0B = 00001011 ( not equal to zero)

According to given rule the LSB of the first octet of the destination address is not equal to zero then it is multicast